# Resonances in b-EMRIs: playing the black hole piano

### v1

@@ II Black hole piano
QNMs are the natural ringing tones of BHs. When a BH is disturbed, it rings at complex QNM frequencies: their real part sets the oscillation frequency (the “note”), while the imaginary part controls the damping rate (how fast the note dies away). The QNM spectrum $_{(s)}\omega_{\ell mn}^{\rm QNM}$ is labeled by the spin-weight $s$ of a perturbation, harmonic numbers $(\ell,m)$ with $\ell\leq m\leq+\ell$ and $\ell\geq s$, and an overtone number $n\geq 0$. Higher overtones are exponentially more damped than the fundamental $n=0$ mode.
@@
where $\tilde{\Omega}=\frac{1}{3\sqrt{3}M}$ denotes the angular velocity of a bound photon orbiting at $\tilde{r}=3M$, while $\gamma_{L}=\frac{1}{3\sqrt{3}M}$ is the Lyapunov exponent governing its orbital instability. In the eikonal limit, the frequency diverges as the total angular momentum, $\omega\sim\ell\to\infty$, while the effects of spin-weight are subleading, so $s$ drops out at order $\mathcal{O}(\ell^{0})$.
@@ III.1 The b-EMRI setup
In this paper, we study the b-EMRI depicted in Fig. 1. The setup consists of an equal-mass binary (hereafter the stellar-mass binary, SB) in the spacetime of a Kerr supermassive BH (SMBH) with mass $M$ and spin parameter $a$. In Boyer-Lindquist coordinates $\{t,r,\theta,\phi\}$, the CoM of the SB is held fixed in the equatorial plane, $\theta=\pi/2$, at a radius $r=r_{0}$. Each component of the SB has mass $\mu$ and executes circular motion about the CoM with separation $d$. We focus on scenarios exhibiting a sharp hierarchy of scales, $\mu\ll d\ll M\lesssim r_{0}$. We allow for a generic inclination of the SB spin with respect to the $\hat{z}-$axis (see Fig. 2), and denote this angle by $\iota_{\rm SB}$.
@@
$\Omega_{\rm SB}=\frac{1}{u^{t}}\sqrt{\frac{2\mu}{d^{3}}}\,$, (2)
@@
where $\mathcal{A}$ depends only on the (both outer and inner) orbital parameters of the SB, $R^{H}\equiv R^{H}_{\ell m\omega_{q}}$ is a solution to the homogeneous Teukolsky equation satisfying purely ingoing boundary conditions at the horizon, and $\bar{S}\equiv{{}_{-2}\bar{S}_{\ell m\omega_{q}}}$. These eigenfunctions are calculated semi-analytically [50, 51, 52, 53, 54]. In a given simulation, for each value of $q$ we must compute the amplitudes $Z_{\ell mq}^{\infty}$ for many harmonics $(\ell,m)$ until we meet a convergence criterion at $\ell=\ell_{\rm max}\sim r_{0}\Omega_{\rm SB}$. This criterion is defined in Eq. (S.5.B) of the Supplemental material of Ref [31].
@@
where the specific expression for $\alpha_{\ell mq}$ can be found in Eq. (4.44) of Ref [55]. The amplitudes on the horizon $Z^{H}_{\ell mq}$ are given by Eq. (5) with $R^{H}\to R^{\infty}$, the solution of the homogeneous Teukolsky equation satisfying outgoing boundary conditions at infinity.
@@ IV.1 An off-center binary in flat space
where $\lambda_{\ell}=\ell(\ell+1)-2$ is the eigenvalue of spin 2-2 spherical harmonics with angular number $\ell$. We call this frequency $\omega^{\rm geom}$ because it is not a special frequency in any way, and the peak in energy flux is entirely due to the geometry of the problem. Recall that in this case the GW signal is just that of a textbook Newtonian binary in the quadrupole approximation; we are simply choosing a bad frame to describe it.
Figure 3: Energy flux in different $\ell$ modes, normalized to total flux $\dot{E}_{Q}$, for a Newtonian binary system with orbital frequency $\Omega_{\rm SB}$ in flat space, with the CoM at $x=r_{0}=10$ (in arbitrary units), $y=z=0$. The spin of the binary points in the $\hat{z}$ direction; we use the quadrupole approximation, so that the GW frequency is simply $2\Omega_{SB}$. Still, the signal is not uniquely described by $\ell=m=2$ harmonics, as would be the case if the CoM were at the origin. Each curve has a global maximum for $2r_{0}\Omega_{\rm SB}\approx 1.1\sqrt{\lambda_{\ell}}$, where $\lambda_{\ell}=\ell(\ell+1)-2$ is the eigenvalue of the spin 2-2 harmonics with angular number $\ell$.
@@
$\frac{Z^{\rm N}_{222}}{\mu d^{2}}=-\frac{\sqrt{5\pi}}{32\,r_{0}^{4}}\eta_{0}^{2}\left(\left(\eta_{0}^{2}-48\right)j_{2}(\eta_{0})+9\eta_{0}\,j_{3}(\eta_{0})\right)\,$, (12)
$\frac{Z^{\rm N}_{332}}{\mu d^{2}}=\frac{\sqrt{42\pi}}{64i\,r_{0}^{4}}\eta_{0}^{2}\left(\left(\eta_{0}^{2}-80\right)j_{3}(\eta_{0})+11\eta_{0}\,j_{4}(\eta_{0})\right)\,$, (13)
@@ IV.2.1 Total energy flux
At low frequencies, $\dot{E}^{Q}$ is not a meaningful estimate of the total flux of the SB, but Fig. 4 still gives us interesting information on how $\dot{E}^{\infty,H}$ behave for small $\Omega_{\rm SB}$ compared to $\dot{E}_{Q}\sim\Omega_{\rm SB}^{6}$. We find the following behavior in the low frequency regime:
@@
This is a counterintuitive result: at low frequencies, the radiation should barely be able to see a BH of size $M$. Nevertheless, most of the radiation is going into the horizon, rather than being scattered to large distances. This behavior has been seen in other setups, and it has been dubbed the “horizon dominance effect” [61, 66]. This is a potentially interesting property of b-EMRIs, as it might lead – when the central BH spins – to overall energy extraction from the BH being deposited into the binary, perhaps stalling the inspiral and leading to floating orbits [67, 68]. We will not explore this further here.
@@ IV.3 Varying the distance from the supermassive BH
Figure 7: Same as Fig. 4 but for different locations $r_{0}\in(2.4M,5M)$ of the SB CoM, with increments of $0.2M$. Varying $r_{0}$ shits the frequency at which the energy fluxes peak, as in the toy model presented in Appendix A. The frequencies for which the energy flux at infinity has a maximum decrease as the SB is moved away from the BH; the opposite is true for the maxima of the energy flux at the horizon.
@@
Figure 8: Top: Total energy flux at infinity; Bottom: Energy flux at infinity for individual $\ell$ modes, normalized to $\dot{E}_{Q}$. Results for a b-EMRI identical to that of Fig. 4 but at $r_{0}=10M$. The peaks in the flux of individual modes do not, in general, correspond to any peak in the total flux. Indeed, the structure of excitation of individual modes is extremely similar to that of flat space (see Fig. 3), where there are no QNMs to excite. We conclude that only the total flux can be used as a smoking gun for excitation of modes of the SMBH.
@@ IV.4 An inclined binary
If the SB emitted isotopically, changing $\iota_{\rm SB}$ would not change the amount of QNM excitation. If it emitted only along the axis, it would be maximal when $\iota_{\rm SB}=\iota_{\rm crit}=\pi/2-\alpha_{\rm crit}$. However, neither is the case; in its local rest frame, the SB emits radiation according to the Peters-Matthews formula [56], so it is not trivial to predict which inclination $\iota_{\rm max}$ maximizes the response of the system. Using an analytical estimate outlined in Appendix C, we find that $\iota_{\rm max}$ has a very sharp transition from $\iota_{\rm max}\approx 0$ when $r_{0}=3M$ to $\iota_{\rm max}\approx\pi/2$ at infinity, with the transition taking place around $r_{0}=4M$.
Figure 10: Total energy flux $\dot{E}=\dot{E}^{\infty}+\dot{E}^{H}$, normalized to $\dot{E}_{Q}$, for a SB at a distance $r_{0}$ from the SMBH and with intrinsic spin at various angles $\iota_{\rm SB}$ relative to the $\hat{z}-$axis (geometric setup in Fig. 9). Top: SB at the light ring $r_{0}=3M$; Bottom: SB at $r_{0}=10M$. In both plots, the flux is larger when the inclination is zero, but approaches the quadrupole formula in the geometric-optics limit for all values of $\iota_{\rm SB}$. When the binary is at the light ring, the resonances (peaks) are most prominent when $\iota_{\rm SB}=0$, as in that case the spin is tangent to the light ring. By contrast, for $r_{0}=10M$, resonances (peaks) are most prominent when $\iota_{\rm SB}=\pi/2$. We find that $\iota_{\rm SB}\approx 0$ is optimal for $r_{0}\lesssim 4$, while $\iota_{\rm SB}\approx\pi/2$ maximizes the resonance when $r_{0}\gtrsim 4$.
Let us then revisit the cases of a static CoM SB at $r_{0}=3M$ and $r_{0}=10M$, but with the intrinsic spin at an inclination $\iota_{\rm SB}\neq 0$. Note that by inclining the binary the reflection symmetry refered to in the last paragraph of Sec. III.1 is broken. In Fig. 10 we show the total flux $\dot{E}=\dot{E}^{\infty}+\dot{E}^{H}$ for varying inclinations $\iota_{\rm SB}$. In the geometric optics limit $\omega\equiv 2\Omega_{\rm SB}\to\infty$, the total flux approaches the quadrupole formula (8), as expected.
@@ IV.5 The effect of the supermassive BH spin
Figure 11: Energy flux at infinity $\dot{E}^{\infty}$, normalized to $\dot{E}_{Q}$, for a static CoM SB at the retrograde light ring in the equatorial plane of a Kerr BH with spin $a=0.9M$. The retrograde light ring is at $r\approx 3.9M$, still outside the ergoregion ($r<2M$). The intrinsic spin of the SB points along the light ring, similar to SB II in Fig. 2. This setup breaks the usual symmetry of the our b-EMRI model under reflections about the equatorial plane, which means that modes with $\omega>0$ ($q=+2$) and $\omega<0$ (q=2q=-2) are independent of each other (see Eq. (4)). The dashed lines show the energy flux only in $\omega>0$ modes (red) and $\omega<0$ modes (blue); combining the two yields the black curve. The positive-(negative-)frequency modes capture the radiation emitted (anti-)parallel to the SB intrinsic spin. Thus, the positive-frequency modes feed the light ring directly, as we saw for Schwarzschild in Fig. 4. In contrast, the negative-frequency modes are emitted mostly in a prograde direction, and hence do not feed the local light ring; this leads to an unpredictable resonance profile, more similar to Fig. 8 but with the added difficulty that the Kerr has a far denser resonance spectrum. The overall result is that the resonances are harder to assign to individual modes when the SMBH is rotating.
@@
Having thoroughly investigated the phenomenon of excitation of QNMs of non-spinning BHs, we now briefly visit the generic case of spinning BHs, described by the Kerr metric. Spinning BHs have longer lived QNMs. This means that the resonance should be sharper and better described by the physics of undamped oscilators (Lorentzian curves, etc). On the other hand, their QNM frequencies now depend on the azimuthal number $m$. This makes their spectrum far denser in the frequency plane, possibly making individual resonances harder to identify.
@@
To test these hypotheses, we place a static CoM SB at the retrograde light ring in the equatorial plane of a Kerr BH with spin $a=0.9M$. The retrograde light ring is at $r\approx 3.9M$, still outside the ergoregion ($r<2M$). We also rotate the SB so that its intrinsic spin points along the light ring, analogously to SB II in Fig. 2, in the retrograde direction (that is, in the same direction as the null geodesics).
@@
In Fig. 11 we look at the energy flux at infinity for this source, as a function of the SB frequency. This setup breaks the reflection symmetry about the $z=0$ plane. Thus, the amplitudes don’t have the symmetry mentioned in the last paragraph of Sec. III.1, and modes with positive and negative frequency display different behaviors. In particular, the former are mostly emitted along the SB intrinsic spin (directly feeding the retrograde light ring as in Sec. IV.2), while the latter are emitted opposite to it. For this reason, the modes with $\omega>0$ (red dashed) show clear resonance with the retrograde light ring modes $m=$-$\ell$, contrary to the modes with $\omega<0$ (blue dashed). The overall flux to infinity (black) is dominated by the negative frequency modes, as the positive frequency modes are more absorbed by the BH; thus, the total flux to infinity does not show a clear resonance with retrograde (or any other) modes. [^3] We don’t show the fluxes on the horizon as these don’t give any new information about the system. Qualitatively, their behavior can be inferred from Fig. 11 and the bottom pannel of Fig. 4.

### v2

# Resonances in binary extreme mass ratio inspirals

@@ II Black hole piano
QNMs are the natural ringing tones of BHs. When a BH is disturbed, it rings at complex QNM frequencies: their real part sets the oscillation frequency (the “note”), while the imaginary part controls the damping rate (how fast the note dies away). The QNM spectrum $_{(s)}\omega_{\ell mn}^{\rm QNM}$ is labeled by the spin-weight $s$ of a perturbation, harmonic numbers $(\ell,m)$ with $\ell\leq m\leq+\ell$ and $\ell\geq|s|$, and an overtone number $n\geq 0$. Higher overtones are exponentially more damped than the fundamental $n=0$ mode.
@@
where $\tilde{\Omega}=\frac{1}{3\sqrt{3}M}$ denotes the angular velocity of a bound photon orbiting at $\tilde{r}=3M$, while $\gamma_{L}=\frac{1}{3\sqrt{3}M}$ is the Lyapunov exponent governing its orbital instability. In the eikonal limit, the frequency diverges as the total angular momentum, $\omega\sim\ell\to\infty$. The effects of spin-weight are $\mathcal{O}(1/\ell)$, so $s$ drops out at this approximation order.
@@ III.1 The b-EMRI setup
In this paper, we study the b-EMRI depicted in Fig. 1. The setup consists of an equal-mass binary (hereafter the stellar-mass binary, SB) in the spacetime of a Kerr supermassive BH (SMBH) with mass $M$ and spin parameter $a$. In Boyer-Lindquist coordinates $\{t,r,\theta,\phi\}$, the CoM of the SB is held fixed in the equatorial plane, $\theta=\pi/2$, at a radius $r=r_{0}$. Each component of the SB has mass $m_{p}$ and executes circular motion about the CoM with separation $d$. We focus on scenarios exhibiting a sharp hierarchy of scales, $m_{p}\ll d\ll M\lesssim r_{0}$. We allow for a generic inclination of the SB spin with respect to the $\hat{z}-$axis (see Fig. 2), and denote this angle by $\iota_{\rm SB}$.
@@
$\Omega_{\rm SB}=\frac{1}{u^{t}}\sqrt{\frac{2m_{p}}{d^{3}}}\,$, (2)
@@
where $\mathcal{A}$ depends only on the (both outer and inner) orbital parameters of the SB, $R^{H}\equiv R^{H}_{\ell m\omega_{q}}$ is a solution to the homogeneous Teukolsky equation satisfying purely ingoing boundary conditions at the horizon, and $\bar{S}\equiv{{}_{-2}\bar{S}_{\ell m\omega_{q}}}$. These eigenfunctions are calculated semi-analytically [50, 51, 52, 53, 54]. In a given simulation, for each value of $q$ we must compute the amplitudes $Z_{\ell mq}^{\infty}$ for many harmonics $(\ell,m)$ until we meet a convergence criterion at $\ell=\ell_{\rm max}\sim r_{0}\Omega_{\rm SB}$. This criterion is defined in Eq. (S.5.B) of the Supplemental material of Ref. [31].
@@
where the specific expression for $\alpha_{\ell mq}$ can be found in Eq. (4.44) of Ref. [55]. The amplitudes on the horizon $Z^{H}_{\ell mq}$ are given by Eq. (5) with $R^{H}\to R^{\infty}$, the solution of the homogeneous Teukolsky equation satisfying outgoing boundary conditions at infinity.
@@ IV.1 An off-center binary in flat space
where $\lambda_{\ell}=\ell(\ell+1)-2$ is the eigenvalue of spin 2-2 spherical harmonics with angular number $\ell$, and the numerical coefficient is found empirically. We call this frequency $\omega^{\rm geom}$ because it is not a special frequency in any way, and the peak in energy flux is entirely due to the geometry of the problem. Recall that in this case the GW signal is just that of a textbook Newtonian binary in the quadrupole approximation; we are simply choosing a bad frame to describe it.
Figure 3: Energy flux in different $\ell$ modes, normalized to total flux $\dot{E}^{Q}$, for a Newtonian binary system with orbital frequency $\Omega_{\rm SB}$ in flat space, with the CoM at $x=r_{0}=10$ (in arbitrary units), $y=z=0$. The spin of the binary points in the $\hat{z}$ direction; we use the quadrupole approximation, so that the GW frequency is simply $2\Omega_{SB}$. Still, the signal is not uniquely described by $\ell=m=2$ harmonics, as would be the case if the CoM were at the origin. Each curve has a global maximum for $2r_{0}\Omega_{\rm SB}\approx 1.1\sqrt{\lambda_{\ell}}$, where $\lambda_{\ell}=\ell(\ell+1)-2$ is the eigenvalue of the spin 2-2 harmonics with angular number $\ell$.
@@
$\frac{Z^{\rm N}_{222}}{m_{p}d^{2}}=-\frac{\sqrt{5\pi}}{32\,r_{0}^{4}}\eta_{0}^{2}\left(\left(\eta_{0}^{2}-48\right)j_{2}(\eta_{0})+9\eta_{0}\,j_{3}(\eta_{0})\right)\,$, (12)
$\frac{Z^{\rm N}_{332}}{m_{p}d^{2}}=\frac{\sqrt{42\pi}}{64i\,r_{0}^{4}}\eta_{0}^{2}\left(\left(\eta_{0}^{2}-80\right)j_{3}(\eta_{0})+11\eta_{0}\,j_{4}(\eta_{0})\right)\,$, (13)
@@ IV.2.1 Total energy flux
At low frequencies, $\dot{E}^{Q}$ is not a meaningful estimate of the total flux of the SB, but Fig. 4 still gives us interesting information on how $\dot{E}^{\infty,H}$ behave for small $\Omega_{\rm SB}$ compared to $\dot{E}^{Q}\sim\Omega_{\rm SB}^{6}$. We find the following behavior in the low frequency regime:
@@
This is a counterintuitive result: at low frequencies, the radiation should barely be able to see a BH of size $M$. Nevertheless, most of the radiation is going into the horizon, rather than being scattered to large distances. This behavior has been seen in other setups, and it has been dubbed the “horizon dominance effect” [61, 66]. This is a potentially interesting property of b-EMRIs, as it might lead – when the central BH spins and can thus power superradiance [67] – to overall energy extraction from the BH being deposited into the binary, perhaps stalling the inspiral and leading to floating orbits [68, 69]. We will not explore this further here.
@@ IV.3 Varying the distance from the supermassive BH
Figure 7: Same as Fig. 4 but for different locations $r_{0}\in(2.4M,5M)$ of the SB CoM, with increments of $0.2M$. Varying $r_{0}$ shifts the frequency at which the energy fluxes peak, as in the toy model presented in Appendix A. The frequencies for which the energy flux at infinity has a maximum decrease as the SB is moved away from the BH; the opposite is true for the maxima of the energy flux at the horizon.
@@
Figure 8: Top: Total energy flux at infinity; Bottom: Energy flux at infinity for individual $\ell$ modes, normalized to $\dot{E}^{Q}$. Results for a b-EMRI identical to that of Fig. 4 but at $r_{0}=10M$. The peaks in the flux of individual modes do not, in general, correspond to any peak in the total flux. Indeed, the structure of excitation of individual modes is extremely similar to that of flat space (see Fig. 3), where there are no QNMs to excite. We conclude that only the total flux can be used as a smoking gun for excitation of modes of the SMBH.
@@ IV.4 An inclined binary
If the SB emitted isotropically, changing $\iota_{\rm SB}$ would not change the amount of QNM excitation. If it emitted only along the axis, it would be maximal when $\iota_{\rm SB}=\iota_{\rm crit}=\pi/2-\alpha_{\rm crit}$. However, neither is the case; in its local rest frame, the SB emits radiation according to the Peters-Matthews formula [56], so it is not trivial to predict which inclination $\iota_{\rm max}$ maximizes the response of the system. Using an analytical estimate outlined in Appendix C, we find that $\iota_{\rm max}$ has a very sharp transition from $\iota_{\rm max}\approx 0$ when $r_{0}=3M$ to $\iota_{\rm max}\approx\pi/2$ at infinity, with the transition taking place around $r_{0}=4M$.
Figure 10: Total energy flux $\dot{E}=\dot{E}^{\infty}+\dot{E}^{H}$, normalized to $\dot{E}^{Q}$, for a SB at a distance $r_{0}$ from the SMBH and with intrinsic spin at various angles $\iota_{\rm SB}$ relative to the $\hat{z}-$axis (geometric setup in Fig. 9). Top: SB at the light ring $r_{0}=3M$; Bottom: SB at $r_{0}=10M$. In both plots, the flux is larger when the inclination is zero, but approaches the quadrupole formula in the geometric-optics limit for all values of $\iota_{\rm SB}$. When the binary is at the light ring, the resonances (peaks) are most prominent when $\iota_{\rm SB}=0$, as in that case the spin is tangent to the light ring. By contrast, for $r_{0}=10M$, resonances (peaks) are most prominent when $\iota_{\rm SB}=\pi/2$. We find that $\iota_{\rm SB}\approx 0$ is optimal for $r_{0}\lesssim 4M$, while $\iota_{\rm SB}\approx\pi/2$ maximizes the resonance when $r_{0}\gtrsim 4M$.
Let us then revisit the cases of a static CoM SB at $r_{0}=3M$ and $r_{0}=10M$, but with the intrinsic spin at an inclination $\iota_{\rm SB}\neq 0$. Note that by inclining the binary the reflection symmetry referred to in the last paragraph of Sec. III.1 is broken. In Fig. 10 we show the total flux $\dot{E}=\dot{E}^{\infty}+\dot{E}^{H}$ for varying inclinations $\iota_{\rm SB}$. In the geometric optics limit $\omega\equiv 2\Omega_{\rm SB}\to\infty$, the total flux approaches the quadrupole formula (8), as expected.
@@ IV.5 The effect of the supermassive BH spin
Figure 11: Energy flux at infinity $\dot{E}^{\infty}$, normalized to $\dot{E}^{Q}$, for a static CoM SB at the retrograde light ring in the equatorial plane of a Kerr BH with spin $a=0.9M$. The retrograde light ring is at $r\approx 3.9M$, still outside the ergoregion ($r<2M$). The intrinsic spin of the SB points along the light ring, similar to SB II in Fig. 2. This setup breaks the usual symmetry of the our b-EMRI model under reflections about the equatorial plane, which means that modes with $\omega>0$ ($q=+2$) and $\omega<0$ (q=2q=-2) are independent of each other (see Eq. (4)). The dashed lines show the energy flux only in $\omega>0$ modes (red) and $\omega<0$ modes (blue); combining the two yields the black curve. The positive-(negative-)frequency modes capture the radiation emitted (anti-)parallel to the SB intrinsic spin. Thus, the positive-frequency modes feed the light ring directly, as we saw for Schwarzschild in Fig. 4. In contrast, the negative-frequency modes are emitted mostly in a prograde direction, and hence do not feed the local light ring; this leads to an unpredictable resonance profile, more similar to Fig. 8 but with the added difficulty that the Kerr has a far denser resonance spectrum. The overall result is that the resonances are harder to assign to individual modes when the SMBH is rotating.
@@
Having thoroughly investigated the phenomenon of excitation of QNMs of non-spinning BHs, we now briefly visit the generic case of spinning BHs, described by the Kerr metric. Spinning BHs have longer lived QNMs. This means that the resonance should be sharper and better described by the physics of undamped oscillators (Lorentzian curves, etc). On the other hand, their QNM frequencies now depend on the azimuthal number $m$. This makes their spectrum far denser in the frequency plane, possibly making individual resonances harder to identify.
@@
To test these hypotheses, we place a stationary CoM SB (at rest with respect to the orbits of $\partial_{t}$) in the retrograde light ring in the equatorial plane of a Kerr BH with spin $a=0.9M$. The retrograde light ring is at $r\approx 3.9M$, still outside the ergoregion ($r<2M$). We also rotate the SB so that its intrinsic spin points along the light ring, analogously to SB II in Fig. 2, in the retrograde direction (that is, in the same direction as the null geodesics).
@@
In Fig. 11 we look at the energy flux at infinity for this source, as a function of the SB frequency. This setup breaks the reflection symmetry about the $z=0$ plane. Thus, the amplitudes don’t have the symmetry mentioned in the last paragraph of Sec. III.1, and modes with positive and negative frequency display different behaviors. In particular, the former are mostly emitted along the SB intrinsic spin (directly feeding the retrograde light ring as in Sec. IV.2), while the latter are emitted opposite to it. For this reason, the modes with $\omega>0$ (red dashed) show clear resonance with the retrograde light ring modes $m=$-$\ell$, contrary to the modes with $\omega<0$ (blue dashed). The overall flux to infinity (black) is dominated by the negative frequency modes, as the positive frequency modes are more absorbed by the BH; thus, the total flux to infinity does not show a clear resonance with retrograde (or any other) modes. [^3] We don’t show the fluxes on the horizon as these don’t give any new information about the system. Qualitatively, their behavior can be inferred from Fig. 11 and the bottom panel of Fig. 4.